AGw Commissioning with IRTC - 11-12 November 2008

Observers: J. Hill (at beginning) and A. Rakich in Tucson plus J. Storm (at end) in Potsdam
Telescope Operator: W. Wack

Summary

Partly to mostly cloudy with poor seeing at the beginning of the night. Things cleared off with somewhat improved seeing for the second half.

We exercised the new (yesterday) build of GCS. It had one instance where it didn't get a WFS image, but otherwise worked OK.

We used a cloudy period to send some presets from the LUCIFER computer. It worked fine, but sluggishness of the downtown interface prevented us from working the system very hard. We did see the automatic M2 offload work correctly with the stepFocus issued by LUCIFER at the end of its preset.

Andrew found that our new focus correction for Z22 corrections with FCin=0 (minimum bending) was incorrect.

Andrew made what should be the final 11 arcsec adjustment of M2 to zero the binodal astigmatism.

Jesper used the last few hours for some measurements on M38, but they had problems with presets from IDL and with the oacserver.

Details (Times reported in MST below)

Open

17:50 Open Dome in Partly Cloudy skies. T=-4 degC

17:52 Home AGw Stages

17:55 pointto M5_0881 gs=0 ACQUIRE

Nov 11 17:55:35 lbtmu05 irc[9499]: MCSPU error returned - ERROR : no tracking polynomials are being received!

17:57 Wayne is able to send a "preset" from the PRESETGUI

17:59 pointto M5_0881 gs=0 ACQUIRE (repeated attempts fail with the same error)

18:02 Stop and restart PCS

18:04 pointto M5_0881 gs=0 ACQUIRE - succeeded but without a star in the frame

18:05 Setting IE=-30 CA=-52 from last night's values. Found star at offset 60 0. New IE=-2.0 CA=-63.0

Exercising new GCS in ACTIVE mode

18:20 pointto M5_0881 gs=3 ACTIVE - Having a slow time converging on the coma because of 2 arcsec seeing.

18:30 ssh'd into lbtmu03 and started the NX interface to the V880 server. It shows the V880 desktop downtown, but hides too much of the downtown desktop.

18:40 In the meantime, GCS has collimated, but the seeing is still 1.4 arcsec.

18:50 appalling seeing, checking Z22 focus compensation. first offload M2 corrections with "focus 0" in IRAF.

18:52 Collimating /tmp wfsc 74 pause wavefront sensing. Cloud rolls over star. Tried going to BS9107 but too low at the moment

19:06 going back to M5_0881 in Acquire mode. Adjust IE=-10 CA=-65

19:10 Resend in Active mode, but WFS images not coming down, GCSL times out waiting for images. see for example 19:19:52. stopped and restarted AGw. Ended up rehoming stages and stopping and restarting GCS (19:28). Confirmed that AGw stages will not move unless they have been homed.

19:50 seeing continues appalling. Looking into Z22 correction. Sent active preset to M5_0881 seeing so bad don't know if it will converge.../tmp wfsc 83-86 for example. Surprisingly WFS converges to ~ 300 nm in absolutely appalling seeing.

20:05 testing sign for M2 coma free rotation adjustment: +x in "Tip" in the WAVEFRONT GUI for M2 is corrected by -X in IE, so IE and "Tip" have a consistent sign convention.

20:20 despite appaling seeing it seems that WFS is converging enough to make tip correction for M2. Resetting IE=-15 CA=-48

20:23 sending guide preset to same star for IDL work on determining direction for M2 tip correction. Preset failed 20:23:56 in GCS log. Clouds keep spoiling our fun.

LUCIFER Preset Tests

20:52 Have Wayne set authorization as LUCIFER.

20:53 ssh into V880 and start the LUCIFER software.

20:56 Stop Mirror Ventilation to reduce load on utility building chiller. Lower DX mirror to reduce compressor load.

21:05 Preset with LUCIFER to FS 106 ACTIVE - Preset timed out after 300 sec, with no hint of the source of the problem in the messages returned to LUCIFER. Based on the indicators on the IIFGUI (left "mount and rotator" is red), I suspect the rotator.

21:25 Wayne revives the rotator, although we don't know why it faulted.

21:28 Preset with LUCIFER to FS 106 ACTIVE. Saw the M2 Offload happen at stepFocus. - Preset seems to have succeed except that there was not any star. GCS reports that it is waiting for the star to be visible. Shouldn't that have made an error in the preset? Maybe the star was briefly seen at acquisition?

21:35 Sent the preset again This time the preset failed, and correctly reported that it could not find a star.

21:42 Preset Telescope to WT10_386 with LUCIFER ACTIVE - Preset failed correctly as there is only a wisp of this bright star on the acquisition image.

21:47 Offset failed since guiding not active.

Back to being IRTC

21:52 pointto BS9107 gs=0 ACQUIRE

21:53 Clear Active Optics and Remove Secondary Offload, Restore Zglobal=0.2 mm

21:55 pointto BS9107 gs=0 ACTIVE collimating on highly aberrated image /tmp wfsc 100-111 and seem to have converged though the Polaris Monitor, Guide image and WFSC image all are in agreement that seeing is >2".

22:07 sending guide preset to GS 18 and checking orientation

22:23 Had an authorization problem with IRS, Doug working on fix (for IDL collimation).

Z22 focus compensation test

23:21 sent -500 nm Z22 expected focus compensation of 0.5 * -22.8*.03789 microns i.e. 443 microns. but in the end the best image turned up with focus compensation of ~ 180 microns. the experiment...

23:30 this time will collimate with the Primary mirror active so don't get focus mixed in with and M2 measurments.

23:45 seeing suddenly much better, ~ 1". Will send 100nm Z22 and note the final position of focus when collimation reconverges. The calculation mentioned above will send 88.6 microns, so I will make a guess that it should be 33.7 microns based on the last measurement and take that out manually before collimating.

00:06 conclusion is that we have the sign and coefficient both wrong for Z22. We are using a factor of -22.8 or so for the ratio of Z4 to Z22, and Z4 goes as 37.89 microns per 1000 nm, so for +100 nm Z22 we are currently sending ~-86.4 microns of delta Z. Looks like we want to send about 055 microns delta Z instead. So, will get this implemented and re-test.

00:25 second test confirmed sign and magnitude wrong, and +100 nm of Z22 wants +53 microns of delta Z on M1 to a close approximation. In the course of doing this have adjusted Z22 down from 147 nm to 4 nm.

M2 Adjustment to Zero the field aberrations

00:30 moving on to M2 adjustment, collimating on-axis on BS9107 (and waiting for around 10-20 minutes for the rotator to slow down as we move down from Zenith).

00:50 collimating with M1 set to active to save confusion. Collimated on axis, now looking for star with Doug's angle around zero or 180 degrees.

01:20 had to ptmodify IE=-2 CA=-51 back to BS9107 collimate on axis, finish at /081112w wfsc104, then preset in guide mode to GS97

01:30 measure GS 97 wfsc 105-107= -376, Z4 = -449, -495, -424

01:34 back on axis, tip M2 by +20 " and adjust IE by -20 " (to -22"). collimating on-axis. wfsc 108-110

01:36 back to GS 97 and re-measuring focus wfsc 111-112 Z4 -512, -648 looks like it went the wrong way. (these are unscaled Z4 by the way, scale up by 1.5) back on-axis to collimate...

01:40 collimating wfsc 113-114 readjusting M2 this time in the opposite direction: Tip -22 " IE +18" collimate wfsc 115-116, preset to 97 again to measure focus Z4 -164, -306, -350 redoing and pushing collimate buttone between measurements -364, -324, -255, back on axis to collimate, looks like that was the right way though.

01:55 readjusted tip to the value of -11" that I think it should be for binodal astigmatism, IE 8, collimating wfsc 123-127, going to GS 80 to measure, no star found, back to on-axis IE 11 CA -60 collimated wfsc -131, going to GS 47

02:06 measuring wfsc 132-134 DA= -238. Y = 485 back on axis to collimate...

02:08 collimating wfsc 135-136 going to GS 82 measuring wfsc 137-139 DA= -290, Y = 485 back on axis to collimate

02:12 collimating wfsc 140-142 going to GS 39 measuring wfsc 143-145 DA = -491 Y = 476.7 back on axis to collimate

02:18 collimating on axis wfsc 146- going to GS 11 measuring wfsc 150-152 DA = -531 y = 440 back on axis etc.

02:24 collimating on axis AGw stopped sending wfsc images, restarting AGw, rehoming stages.

02:36 back in action collimated to wfsc 156 going to GS 20 measuring wfsc 157-159, DA = -201 y = 442, back on axis (BOA)

02:38 collimating wfsc 160-167 seeing worse, going to GS 96 measuring wfsc 168-170 DA = -329 y = 457, BOA

02:50 collimating wfsc 171-174 , going to GS 106 measuring wfsc 175-177, DA = -361 y = 455, BOA

02:56 collimating wfsc 178-181 , going to GS 8 measuring wfsc 182-184, DA=-185 , y =432, BOA

03:03 collimating 185- seeing degraded below usefulness keeping updating collimation till it passes...getting cores back in the wfsc images (wfsc 194)

03:14 going to GS 62, measuring wfsc 197-199, DA = -453, y = 478.5 probably a little faint to be useful 250 counts max on wfsc, BOA,

03:18 collimating wfsc 200-201, going to GS 64 measuring wfsc 202-204, DA = -452 , y = 446,

03:26 back on axis collimating wfsc 204-207, then hit -y limit on M1 added +1 mm to Y on M1 and M2 BOA,

03:30 collimating wfsc 207- 210 going to GS 10, measuring wfsc 211-213, DA = -204 , y = 428, BOA

03:35 collimating wfsc 214-220, going to GS 35 measuring wfsc 221-223, DA = -231, y = 464, BOA,

03:44 collimating wfsc 224-230, going to GS 70 measuring wfsc 231-233, DA = -284, y =457 , now handing over to Jesper for remainder of night.

AGw Calibrations

03:53 going on-axis in acquire mode to tweak IE CA if necessary-> not necessary IE=+11 CA=-60 with new M2 position

04:30 collimating on a mag 10 star nearby to Jespers target M38.

05:03 after sorting out numerous things with a woken up Doug are trying again. Having trouble with newly conceived "track" and "guide" IRC presets.

05:12 AGw stage runs to end of its limits in Y and Z for no known reason, suspect the newly written irc_preset commands in IDL (given that they were put together at 4:00 am by someone who had just woken up).

05:16 AGw refusing to home, trying restart of OACserver

05:20 OACserver failing to restart

05:33 must be serious... Thomas Hahn called in

05:45 got OAC server started up again after finding typo in new config file. Got data for Jesper on M38 going to M67.

05:55 for some unknown reason the irc_preset_guide preset is setting focus to zero, see for example the GCS log at 05:58:31, given up on getting a last set of data before sunrise, closing up.

Ended the night on IE=+11 CA -60

-- JohnHill - 12 Nov 200
Topic revision: r9 - 12 Nov 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback